# Metasurface-based Snapshot Shortwave-Infrared Hyperspectral Image Reconstruction with Inter and Intra Prior Learning Network

### v2

@@ IV Experiments
In this section, we compare our method with several state-of-the-art(SOTA) methods on the AVIRIS-NG(Airborne Visible InfraRed Imaging Spectrometer Next Generation) dataset. We selected 300 wavelengths ranging from 1000 to 2500 and removed the bands whose response values were zero due to water vapor absorption. The peak-signal-to-noise-ratio (PSNR) and structured similarity index metrics (SSIM) metrics are used to evaluate the performance of different hyperspectral image reconstruction methods.
@@ IV-A Datasets
AVIRIS-NG measures the wavelengths ranging from 380 to 2510 nm with a 5 sampling interval. Spectra are captured as images with 600 cross-track elements and spatial sampling ranging from 0.3 m to 4.0 m. We selected 12 flight-lines, and each hyperspectral image is approximately sized at . We create spatial domain tiles for each image, with each tile sized at .
@@ IV-B Implementation Details
Our model was implemented using the PyTorch framework and trained with the Adam optimizer for epochs. During the training, we applied randomly cropping, rotation, and flipping of tiles for data augmentation. The learning rate was set to and the batch size 1. The training and test data size was . All experiments were conducted on the NVIDIA Tesla P40 GPU.
@@ IV-C Evaluation Indexes
We adopt four image quality indexes, including peak signal-to-noise ratio (PSNR) and structure similarity (SSIM)[67], for quantitative evaluation. Specifically, PSNR measures the visual quality, SSIM measures the structure similarity. Generally, bigger values of PSNR and SSIM suggest a better reconstruction accuracy.
@@ IV-D Comparisons with SOTA Methods
We conducted a comprehensive comparative analysis of the proposed method and various SOTA methods, including two end-to-end methods(TSANet,HDNet) and three deep unfolding networks(GAP-Net,ADMM-Net,RDLUF). It is worth noting that RDLUF is implemented with help of the cross-stages interaction and multi-scales encoder feature fusion. We replaced the CASSI simulation imaging part of all comparison methods with the simulation process of filter array imaging, adjusting the number of feature channels, and maintaining consistency in learning rate, loss function, optimizer settings as much as possible.
@@ IV-E Ablation Study
Break-down Ablaton. We used a plain ISTA gradient descent module followed by a spatial-spectral prior learning(SSPL)-based U-Net as our baseline model. Subsequently, incorporating low-rank prior learning(LRPL) block to better extract global low rank prior information led to an improvement of . Adding a adaptive feature transfer(AFT) block to reduce feature loss between different scales, achieving an additional improvement of .
Number of stages. To explore the impact of the number of stages, we analyze the variation of model performance ranging from 3 to 9 stages as presented in TABLE I. We find that the reconstruction quality improves with the increase in the number of stages, indicating the high efficiency of the iterative deep unfolding network design. However, the reconstuction quality improve slow down after nine stages. After weighing the balance between computational complexity and performance improvement, we set the number of stages to nine.
| SSPL | LRPL | AFT | |
| PSNR | 40.46 | 41.13 | 41.57 |
| SSIM | 0.984 | 0.988 | 0.989 |
| Paras | 390.41M | 455.57M | 517.46M |
| GFLOPs | 1124.81 | 1458.77 | 1561.29 |
@@
Ablation of Metasurface Optimization. In order to verify the effectiveness of the our metasurface optimization method combined with the farthest distance sampling and Pearson correlation coefficient, we conduct experiments under the filter array selected by two methods(naive minimize person correlation coefficient method and our method) as shown in table 4. The experimental result indicates that compared with considering the minimum correlation on the whole metasurface database, our proposed method shows great superiority in HSI reconstruction quality. Besides, the visualization results also show that the randomness of the metasurface transmittance curve optimized by our method better than original method, and the correlation between the metasurface units in the filter array changes more smoothly, we believe that the smoothness can improve the stability of the reconstruction network training process, further improve reconstruction performance.
@@ V Conclusion
In this paper, we enhance the Pearson correlation coefficient based filter selection and propose a inter and intra prior learning deep unfolding network for SWIR-HSI reconstruction. By intefrating FPS into the original filter selection, we achieve better optimization of filter arrays and higher reconstruction quality. To address the loss of detailed features in the network, a adaptive feature transfer module is introduced to adaptively fuse the detail information from encoder stages and transfer them to decoder stages. Experimental results demonstrate that our method outperforms previous HSI reconstruction network.
Limitations and social impact. The proposed framework extends spectral filter array imager to SWIR wavelengths and contributes to the design of deep unfolding networks for other inverse problems. However, our framework currently cannot handle large-scale NIR-HSI due to the high number of bands involved.

### v3

@@ IV Experiments
In this section, we compare our method with several state-of-the-art (SOTA) methods on the AVIRIS-NG (Airborne Visible InfraRed Imaging Spectrometer Next Generation) dataset. We selected 300 wavelengths ranging from 1000nm to 2500nm and removed ands whose response values were zero due to water vapor absorption. The peak-signal-to-noise-ratio (PSNR) and structured similarity index (SSIM) metrics are used to evaluate the performance of different hyperspectral image reconstruction methods.
@@ IV-A Datasets
AVIRIS-NG measures wavelengths ranging from 380nm to 2510nm with a 5nm sampling interval. Spectra are captured as images with 600 cross-track elements and spatial sampling ranging from 0.3m to 4.0m. We selected 12 flight lines, and each hyperspectral image is approximately sized at 21K 0.65K 432. We create spatial domain tiles for each image, with each tile sized at 256 256 300.
@@ IV-B Implementation Details
Our model was implemented using the PyTorch framework and trained with the Adam optimizer for epochs. During the training, we applied random cropping, rotation, and flipping of tiles for data augmentation. The learning rate was set to 0.0002 and the batch size 1. The training and test data size was 64 64 300. All experiments were conducted on the NVIDIA Tesla P40 GPU.
@@ IV-C Evaluation Indexes
We adopt two image quality indexes, including peak signal-to-noise ratio (PSNR) and structure similarity (SSIM)[66], for quantitative evaluation. Specifically, PSNR measures the visual quality, SSIM measures the structure similarity. Generally, bigger values of PSNR and SSIM suggest a better performance.
@@ IV-D Comparisons with SOTA Methods
We conducted a comprehensive comparative analysis of the proposed method and various SOTA methods, including two end-to-end methods (TSANet[21],HDNet[9]) and three deep unfolding networks (GAP-Net[29],ADMM-Net[28],RDLUF[30]). It is worth noting that RDLUF is implemented with help of the cross-stages interaction and multi-scales encoder feature fusion. We replaced the CASSI simulation imaging part of all comparison methods with the simulation process of filter array imaging, adjusting the number of feature channels, and maintaining consistency in learning rate, loss function, optimizer settings as much as possible.
@@
To compare our method with other state-of-the-art methods qualitatively, we have selected several methods and showcased the results in Scene03 across five spectral bands in Fig. 4. The reconstruction results of our method and five others are presented, along with the respective residual maps for each reconstruction. Notably, our method demonstrates lower reconstruction errors, especially in the top right corner. Furthermore, in a grassland texture-based Scene05, as depicted in Fig. 5, our method closely aligns with the ground truth in terms of visual effect, while other methods exhibit fewer details and more artifacts in fine features.
| Method | Params(M) | FLOPs(G) | PSNR(dB) | SSIM |
| TSANet | 162.87 | 361.37 | 39.74 | 0.984 |
| ADMM-Net | 266.37 | 331.24 | 40.03 | 0.985 |
| GAP-Net | 266.37 | 331.24 | 40.06 | 0.985 |
| HDNet | 529.64 | 2167.27 | 41.05 | 0.988 |
| RDLUF | 289.63 | 974.577 | 41.33 | 0.988 |
| Ours-9stag | 517.46 | 1561.29 | 41.57 | 0.989 |
For a more detailed quantitative analysis of the reconstructions in local regions, we illustrated two selected regions from Scene03 and Scene05 in Fig. 6, labeled as a, b, c, and d. We presented the spectral curves and computed their correlation with the ground truth data, comparing the results with five state-of-the-art methods. Our method’s spectral curve is represented by the black curve, closely resembling the red curve of the ground truth data.
In a comprehensive comparison, we analyzed the complexity of our proposed algorithm alongside other state-of-the-art reconstruction algorithms, as displayed in Table II. While our method demonstrates a moderate level of computational complexity, it requires significantly more floating-point operations compared to the deep unfolding method RDLUF [30].
@@ IV-E Ablation Study
Break-down Ablaton. To assess the efficacy of the different components in our approach, we conducted a series of ablation experiments with k set to 9 stages. Table III presents the PSNR and SSIM results for each component. Initially, we substituted the complete ERRA network with a spatial-spectral prior learning architecture as the base framework, as shown in row (a) of Table III. The LRPL module demonstrated an increase of 0.67dB in PSNR compared to the base framework, showcasing the effectiveness of our cross-stage low rank prior learning method, as seen in row (b) of Table III. Additionally, we integrated the Adaptive Feature Transfer (AFT) module into the LRPL framework to assess the interaction of feature information. The results in row (c) of Table III indicate that the AFT module achieved a 0.44dB higher PSNR compared to the LRPL framework, confirming its efficacy across different stages of the network.
Number of stages. We conducted experiments at various stages to analyze the impact of the unfolding network on overall expressiveness. The results are presented in Table I, showcasing the computational complexity and average reconstruction quality on the AVIRIS-NG test dataset. With an increasing number of network stages, there is a gradual rise in floating-point calculations, resulting in improved experimental outcomes. To find a balance between impact and complexity, we settled on 9 as the final number of stages for the network.
| setting | SSPL | LRPL | AFT | PSNR(dB) | SSIM |
| a | ✓ | 40.46 | 0.984 | ||
| b | ✓ | ✓ | 41.13 | 0.988 | |
| c | ✓ | ✓ | ✓ | 41.57 | 0.989 |
Moreover, as the network progresses through more stages, there are minimal parameter changes due to our strategy of parameter sharing across stages during training. Additionally, a line graph in Fig. 8 visually represents the stage changes and the network’s evolution. The horizontal axis denotes the number of stages (k) in the unfolding network, the left axis shows the PSNR results, and the right axis displays the SSIM values. Notably, as the stage number increases, both PSNR and SSIM metrics show progressive enhancement.
IV-F Pattern layout analysis
To verify the importance of the pattern layout in enhancing the reconstruction quality, we conducted a comparative experiment using different filter selection methods and varying filter numbers.
Number of metasurface filter. For the optimization of the basic filter array, we replaced the basic filter array with different number of filters(i.e, Case 1, Case 2, Case 3) as shown in Fig. 7, and all of the basic filter arrays were generated using our proposed filter selection method. The experimental results indicate that a smaller number of filters leads to better reconstruction performance as presented in Table IV. Since the spectral information required for reconstructing the spectrum of each pixel is captured by various filters within the local region, a large number of filters in basic array may introduce incorrect spectral information that differs from the central pixel due to the significant variations in spectral features of remote sensing images. This can adversely affect spectral reconstruction and result in a decrease in reconstruction performance. Hence, we have chosen Case 3 as our standard basic filter array.
| Metric | Case 1 | Case 2 | Case 3 | Case 4 |
| PSNR(dB) | 40.17 | 40.41 | 41.20 | 40.57 |
| SSIM | 0.984 | 0.985 | 0.988 | 0.983 |
@@ IV-F Pattern layout analysis
Filter selection method. In order to verify the effectiveness of the our proposed filter selection method combined with the farthest distance sampling and Pearson correlation coefficient, we conducted experiments under the filter array selected by two methods (inner product-based optimization method[14] and our method), the transmissivity of existing method is shown in Fig. 9(a)(Case 4). As presented in TableIV, existing filter selection method(Case 4) shows a large lower than our method(Case 3) by 0.63dB in PSNR and 0.005 in SSIM. Besides, the visualization results also show that the randomness of metasurface filter selected by our method better than existing method, and the correlation between the metasurface units in the basic filter array changes more smoothly as shown in Fig. 9(b-c).
@@ V Conclusion
In this paper, we enhance the correlation coefficient based filter selection and propose an inter and intra prior learning deep unfolding network for SWIR-HSI reconstruction. By intefrating FPS into the original filter selection, we achieve better optimization of filter arrays and higher reconstruction quality. For integrating low rank prior learning and cross-stage information interaction, a cross-stage low rank prior learning module is proposed to globally capture HSI’s useful characteristics. To address the loss of detailed features in the network, a adaptive feature transfer module is introduced to adaptively fuse the detail information from encoder stages and transfer them to decoder stages. Experimental results demonstrate that our method outperforms previous HSI reconstruction network.